OpenStudy (anonymous):

how much should be invested at 10% if one has $5000 to invest, in order that the annual income from one investment at 10% shall double the income from another investment at 15%?

OpenStudy (phi):

They don't say, but let's assume this is simple interest, so the annual income is the interest: Interest Rate * Principal they question seems to be asking: if you divide 5000 into 2 parts (call them x and y) then the interest on the x part (at 10%) is twice as big as the interest on the y part (invested at 15%) here is what you have: x+y= 5000 (the sum of the two chunks of money is 5000) 10% of x = 2*(15% of y) can you solve this ?

OpenStudy (anonymous):

I dont get it :(

OpenStudy (phi):

what is the interest you get from 5000 at 10% ?

OpenStudy (anonymous):

500

OpenStudy (phi):

yes 0.10*5000=500 but if you put all of the 5000 in at 10% there is none left over to invest at 15% the interest you get on the amount invested at 15% is 0 (because we invested nothing) if we invested half the money at 10% and half at 15% we would get 0.10*2500 = 250 0.15*2500 = 375 but the question says we want ** one investment at 10% shall double the income from another investment at 15%?** in other words they want the interest from 0.10 more (twice as much) as the interest from the 0.15

OpenStudy (phi):

you could try "trial and error" and keep putting more and more into 10% and the amount left over into 15%, and compare the amount of interest. for example 0.10*4000= 400 0.15*1000= 150 now 400 is more than twice as much as the 150, so 4000 is too much to put in the 10%

OpenStudy (phi):

but you can use algebra to solve this: let x be the amount to put into 10% how much is left over for the 15% ? we start with 5000, take away x, what is left ? 5000-x so we could say 0.10*x = some amount = A 0.15*(5000-x) = 1/2 as much = A/2 (call the amount of interest A) if you double the interest from the 15% 2* 0.15*(5000-x) = 2*A/2 = A it will match the interest from 10% set them equal: 2* 0.15*(5000-x) = 0.10*x now simplify and solve for x